the length of a rectangle is 1 cm more than twice the width. the perimeter of the rectangle is 110 cm. find the lenght and the width of the rectangle

Answer by unlockmath(1688)

Hello,
We can set up the formula:
2L +2W = 110
We can also set up length as 2x+1 and width as x so plug this in:
2(2x+1)+2x=110
Expanded to:
4x+2+2x=110
Rewritten as:
6x=108
Divide both sides by 6:
x=18
So width is 18 cm
and length is 37 cm
Make sense?
RJ
www.math-unlock.com